# Wiman-Valiron method for fractional derivatives and sharp growth estimates of $\alpha$-analytic solutions for linear fractional differential equations

### v1

@@ Abstract
We consider a fractional linear differential equation with successive derivatives given by $\mathbb{D}_{\alpha}^{n}y+p_{n-1}(x)\mathbb{D}_{\alpha}^{n-1}y+\dots+p_{1}(x)\mathbb{D}_{\alpha}y+p_{0}(x)y=0$, where $\mathbb{D}_{\alpha}^{j}$ is the $j$th iteration of the Caputo-Djrbashian fractional derivative of order $\alpha>0$, $p_{j}$ are $\alpha$-analytic functions for $0<x^{\alpha}<R$. Generalizing a result of Kilbas, Rivero Rodríguez-Germá and Trujillo, we prove the existence and uniqueness of the corresponding Cauchy problem in the class of $\alpha$-analytic functions. We establish an exact growth order for the solution when $p_{j}(x)=P_{j}(x^{\alpha})$, where $P_{j}$ are polynomials, and $p_{0}$ dominates in some sense. This is the full counterpart of the classical case of ordinary differential equations. In particular, we demonstrate the sharpness of Kochubei’s result and generalize it. To achieve this, we extend the Wiman-Valiron theory to analytic functions and the Djrbashian-Gelfond-Leontiev generalized fractional derivatives.
Keywords: fractional calculus (primary); fractional linear differential equations; Wiman-Valiron method; $\alpha$-analytic solution; Cauchy problem; Caputo-Djrbashian derivative; Gelfond-Leontiev derivative; growth of solutions; Mittag-Leffler functions
@@ 1.1 Fractional linear differential equations
$\mathbb{D}_{\alpha}^{n}y+p_{n-1}(x)\mathbb{D}_{\alpha}^{n-1}y+\dots+p_{1}(x)\mathbb{D}_{\alpha}y+p_{0}(x)y=0,$ (1)
where $\mathbb{D}_{\alpha}$ is the fractional Caputo-Djrbashian derivative, $\mathbb{D}_{\alpha}^{j}$ its $j$th iteration, $p_{j}$, $j\in\{0,\dots,n-1\}$ are $\alpha$-analytic functions, i. e. $p_{j}(x)=P_{j}(x^{\alpha})$, $P_{j}(z)$ being analytic on $\{z\in\mathbb{C}:|z|<R\}$, $0<R\leq\infty$. This type of equation was introduced and studied in the case of constant coefficients and the Riemann-Liouville operator $D^{\alpha}$ instead of the Caputo-Djrbashian operator in [19, Chap. V]. It is worth remarking that neither the semigroup property $D^{\alpha}D^{\beta}=D^{\alpha+\beta}$ nor the commutativity property $D^{\alpha}D^{\beta}=D^{\beta}D^{\alpha}$ holds in general for either the Riemann-Liouville or the Caputo-Djrbashian operators (see [19, Chap. IV.6], [2]). Suffiecient conditions for the semigroup property and commutativity can be found in [19, Chap. IV.6], [1].
Note that the existence and uniqueness of a solution to the Cauchy problem for a more general than (1) linear differential equation of fractional order was established in [9].
As it is stated in [13] for the cases $n=1$ and $n=2$ equation (1) possesses the unique $\alpha$-analytic soultions on $0<|x|^{\alpha}<R$ that satisfy initial conditions for both the Riemann-Liouville and the Caputo-Djrbashian derivatives. Although the formal series representations of solutions are given in [13], their convergence is not rigorously proved. On the other hand, A. Kochubei ([17]) established an asymptotically sharp estimate of the growth of solutions to equation (1) in the case $n=1$ where $A$ is a polynomial. In other words, the Cauchy problem
@@
On one hand, there is still no understanding of how one can generalize the logarithmic derivative estimate approach for fractional derivatives (cf. [6]). On the other hand, in [7] the authors succeeded in generalizing the Wiman-Valiron method for Riemann-Liouville derivatives. This allowed us to obtain sharp asymptotic growth of solutions to a special fractional linear differential equation, but not for (1). The reason is that, to find an asymptotic for an $\alpha$-analytic solution of (1), one needs to generalize the Wiman-Valiron theory for Gelfond-Leontiev type derivatives (see [15], [16]). We do this in Section 3, where we prove Theorem 1, the main result of the paper. In the final section we study equation (1). Theorem 2 establishes the existence and uniqueness of a solution to the Cauchy problem for (1) in the class of $\alpha$-analytic functions. This is a slight generalization of a result from [13]. We then consider the case when all coefficients of (1) are polynomials of $t^{\alpha}$. We prove (Theorem 3) that in this case all $\alpha$-analytic solutions are of the form $v(t^{\alpha})$ where $v$ is an entire function of finite order of the growth. Finally, Theorem 4 gives sharp values for the order of the growth of $v$ under natural conditions on the coefficients. Auxiliary results are given in Section 2.
We use the notation $a\lesssim b$ if there exists a constant $C>0$ such that $a\leq Cb$. Similarly, $a\gtrsim b$ is understood in an analogous manner. If $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$ and say that $a$ and $b$ are comparable. Additionally, $a(t)\sim b(t)$ means that the quotient $a(t)/b(t)$ approaches one as $t$ tends its limit.
@@ 1.2 Fractional integrals and derivatives
$I^{\alpha}x^{\beta-1}=\frac{\Gamma(\beta)}{\Gamma(\beta+\alpha)}x^{\beta+\alpha-1},\;D^{\alpha}x^{\beta-1}=\frac{\Gamma(\beta)}{\Gamma(\beta-\alpha)}x^{\beta-\alpha-1},\quad\alpha,\beta>0,\alpha\neq\beta,$ (2)
$D^{\alpha}1=\frac{1}{\Gamma(1-\alpha)}x^{-\alpha},\;D^{\alpha}x^{\alpha-j}=0,\quad\alpha>0,j\in\{1,2,\dots,[\alpha]+1\}.$ (3)
@@
$=D^{\alpha}\varphi(x)-\sum_{k=0}^{n-1}\frac{\varphi^{(k)}(0)}{\Gamma(k+1-\alpha)}x^{k-\alpha},\quad n-1<\alpha\leq n,$ (4)
@@
Let $H_{\alpha}(R)$, $0<R\leq\infty$, $\alpha>0$, denote the class of $\alpha$-analytic functions on $(0,R^{\alpha})$, that is the functions $u$ represented in the form $u(t)=\sum_{m=0}^{\infty}a_{m}t^{\alpha m}$, $\alpha>0$, $0\leq t^{\alpha}<R\leq\infty$. Direct computation yields
$(D^{\alpha}u)(t)=\sum_{m=0}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-1)\alpha+1)}t^{\alpha(m-1)},\quad 0<t^{\alpha}<R,$ (5)
@@
$=\sum_{m=1}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-1)\alpha+1)}t^{\alpha(m-1)}.$ (6)
@@
$(\mathbb{D}_{\alpha}^{j}u)(t)=\sum_{m=j}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-j)\alpha+1)}t^{\alpha(m-j)}=\sum_{m=0}^{\infty}u_{m+j}\frac{\Gamma((m+j)\alpha+1)}{\Gamma(m\alpha+1)}t^{\alpha m}.$ (7)
@@
If $u\in H_{\alpha}(R)$, then $D^{\alpha}$ is associative and commutative providied the conditions $u_{m}(0)=0$ for $m<\frac{\gamma+\beta}{2}$, $\gamma,\beta>0$, i.e.
@@
In particular, $\left(D^{\alpha}\right)^{j}=D^{j\alpha}$ if $u_{m}(0)=0$ for $m<j$.
@@
We write $u(t)=\sum_{m=0}^{\infty}a_{m}t^{\alpha m}$, $\alpha>0$, $0<t^{\alpha}<R\leq\infty$. Since the power series is uniformly convergent on every segment $0,r^{\alpha}]\subset[0,R^{\alpha})$ (cf. [19, Theorem 3, Sec. IV.6]), we can integrate and differentiate it under the sum sign at every point of $0,R^{\alpha})$.Then using the fact that $u_{m}(0)=0$ for $m<(\gamma+\beta)/\alpha$ we obtain
@@
$=\sum_{m=\left[\frac{\gamma+\beta}{\alpha}\right]}^{\infty}u_{m}\frac{\Gamma(\alpha m+1)}{\Gamma(\alpha m+1-\gamma)}\frac{\Gamma(\alpha m-\gamma+1)}{\Gamma(\alpha m+1-\gamma-\beta)}t^{\alpha m-\gamma-\beta}=D^{\gamma+\beta}u(t).$ (8)
@@
$f(z)=\sum\limits_{n=0}^{\infty}a_{n}z^{n},\quad z=re^{i\theta}$ (9)
@@
$(\mathbb{D}^{\alpha}u)(t)=\sum_{n=1}^{\infty}a_{n}\frac{\Gamma(n\alpha+1)}{\Gamma(n\alpha+1-\alpha)}t^{\alpha(n-1)}=(\mathcal{D}^{\alpha}f)(t^{\alpha}),$ (10)
@@
is the so-called Djrbashian-Gelfond-Leontiev operator ([15], [16]), a special case of the Gelfond-Leontiev generalized differential operator corresponding to the Mittag-Leffler function $E_{\alpha}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(k\alpha+1)}$, $\alpha>0$.
The corresponding integral operator, right inverse to $\mathcal{D}^{\alpha}$, can be written as
$(\mathcal{I}^{\alpha}f)(z)=\sum_{n=0}^{\infty}a_{n}\frac{\Gamma(n\alpha+1)}{\Gamma(n\alpha+1-\alpha)}z^{n+1},$
@@
$\mathcal{I}^{\alpha}f(z)=\frac{z}{\Gamma(\alpha)}\int_{0}^{1}(1-t)^{\alpha-1}f(zt^{\alpha})\,dt.$
It follows from (10) that ([14, Section 2.5], [20, Sec. 18.2, 22.3])
$\mathcal{D}^{\alpha}=Q\circ\mathbb{D}_{\alpha}\circ Q^{-1},$
where $Q$ is the substitution operator $z\mapsto z^{\frac{1}{\alpha}}$, $Q^{-1}$ is its inverse, for a corresponding branch of a multivalued power function chosen on a segment $[0,z]$. This note allows us to apply the approach used in [7] to derive Wiman-Valiron type results for the operator $\mathcal{D}^{\alpha}$.
@@ 1.3 Wiman-Valiron theory
For $r\in[0,+\infty)$ and an entire function $f$ of the form (9) we denote $M(r,f)=\max\{|f(z)|:|z|=r\}$. We define the maximal term as $\mu(r,f)=\max\{|a_{n}|r^{n}:n\geq 0\}$ and the central index of the series as $\nu(r,f)=\max\{n\geq 0:|a_{n}|r^{n}=\mu(r,f)\}$. Note that $\nu(r,f)$ is non-decreasing, and an entire function $f$ is transcendental if and only if $\nu(r,f)\to+\infty$ as $r\to+\infty$.
@@
Let $V$ be the class of positive continuous nondecreasing functions $v$ on $0,+\infty)$ such that $\frac{x^{2}}{v(x)\ln v(x)}$ increases to $\infty$ on $x\in[x_{0};+\infty)$, $x_{0}>0$, and $\int\limits_{0}^{+\infty}\frac{dx}{v(x)}<+\infty$. For example, the function $v(x)=x\ln^{\alpha+1}x$, $(x\geq e)$, $\alpha\in(0,1)$ belongs to $V$.
@@
$\mathcal{D}_{\alpha}^{j}f(z)=(\nu\alpha)^{j\alpha}\frac{f(z)}{z^{j}}+O\left(\frac{\varkappa(\nu)}{\nu}\right)\frac{M(\rho,f)}{\rho^{j}}.$ (12)
@@
$M(\rho,\mathcal{D}_{\alpha}^{j}f(z))=\frac{(\nu\alpha)^{j\alpha}}{\rho^{j}}\left\{1+O\left(\frac{\varkappa(\nu)}{\nu}\right)\right\}M(\rho,f)=(1+o(1))\frac{(\nu\alpha)^{j\alpha}}{r^{j}}M(r,f)$ (13)

### v2

@@ Abstract
We consider a fractional linear differential equation with successive derivatives given by $\mathbb{D}_{\alpha}^{n}y+p_{n-1}(x)\mathbb{D}_{\alpha}^{n-1}y+\dots+p_{1}(x)\mathbb{D}_{\alpha}y+p_{0}(x)y=0$, where $\mathbb{D}_{\alpha}^{j}$ is the $j$th iteration of the Caputo-Djrbashian fractional derivative of order $\alpha>0$, $p_{j}$ are $\alpha$-analytic functions for $0<x^{\alpha}<R$. Generalizing a result of Kilbas, Rivero, Rodríguez-Germá and Trujillo, we prove the existence and uniqueness of the corresponding Cauchy problem in the class of $\alpha$-analytic functions. We establish an exact growth order for the solution when $p_{j}(x)=P_{j}(x^{\alpha})$, where $P_{j}$ are polynomials, and $p_{0}$ dominates in some sense. This is the full counterpart of the classical case of ordinary differential equations. In particular, we demonstrate the sharpness of Kochubei’s result and generalize it. To achieve this, we extend the Wiman-Valiron theory to analytic functions and the Djrbashian-Gelfond-Leontiev generalized fractional derivatives.
Keywords: fractional calculus (primary); fractional linear differential equations; Wiman-Valiron method; $\alpha$-analytic solution; Cauchy problem; Caputo-Djrbashian derivative; Gelfond-Leontiev derivative; growth of solutions; Mittag-Leffler functions; Bell polynomials.
@@ 1.1 Fractional linear differential equations
$\mathbb{D}_{\alpha}^{n}y+p_{n-1}(x)\mathbb{D}_{\alpha}^{n-1}y+\dots+p_{1}(x)\mathbb{D}_{\alpha}y+p_{0}(x)y=0,$ (1.1)
where $\mathbb{D}_{\alpha}$ is the fractional Caputo-Djrbashian derivative, $\mathbb{D}_{\alpha}^{j}$ its $j$th iteration, $p_{j}$, $j\in\{0,\dots,n-1\}$ are $\alpha$-analytic functions, i. e. $p_{j}(x)=P_{j}(x^{\alpha})$, $P_{j}(z)$ being analytic on $\{z\in\mathbb{C}:|z|<R\}$, $0<R\leq\infty$. This type of equation was introduced and studied in the case of constant coefficients and the Riemann-Liouville operator $D^{\alpha}$ instead of the Caputo-Djrbashian operator in [19, Chap. V]. It is worth remarking that neither the semigroup property $D^{\alpha}D^{\beta}=D^{\alpha+\beta}$ nor the commutativity property $D^{\alpha}D^{\beta}=D^{\beta}D^{\alpha}$ holds in general for either the Riemann-Liouville or the Caputo-Djrbashian operators (see [19, Chap. IV.6], [2]). In particular, $\mathbb{D}_{\alpha}^{j}\neq\mathbb{D}_{\alpha j}$. Suffiecient conditions for the semigroup property and commutativity can be found in [19, Chap. IV.6], [1].
Note that the existence and uniqueness of a solution to the Cauchy problem for a more general than (1.1) linear differential equation of fractional order was established in [9].
As it is stated in [13] for the cases $n=1$ and $n=2$ equation (1.1) possesses the unique $\alpha$-analytic soultion on $0<|x|^{\alpha}<R$ that satisfy initial conditions for both the Riemann-Liouville and the Caputo-Djrbashian derivatives. Although the formal series representations of solutions are given in [13], their convergence is not rigorously proved. On the other hand, A. Kochubei ([17]) established an asymptotically sharp estimate of the growth of solutions to equation (1.1) in the case $n=1$ where $A$ is a polynomial. In other words, the Cauchy problem
@@
On one hand, there is still no understanding of how one can generalize the logarithmic derivative estimate approach for fractional derivatives (cf. [6]). On the other hand, in [7] the authors succeeded in generalizing the Wiman-Valiron method for Riemann-Liouville derivatives. This allowed us to obtain sharp asymptotic growth of solutions to a special fractional linear differential equation, but not for (1.1). The reason is that, to find an asymptotic for an $\alpha$-analytic solution of (1.1), one needs to generalize the Wiman-Valiron theory for Gelfond-Leontiev type derivatives (see [15], [16]). We do this in Section 3, where we prove Theorem 1, the main result of the paper. In the final section we study equation (1.1). Theorem 2 establishes the existence and uniqueness of a solution to the Cauchy problem for (1.1) in the class of $\alpha$-analytic functions. This is a slight generalization of a result from [13]. We then consider the case when all coefficients of (1.1) are polynomials of $t^{\alpha}$. We prove (Theorem 3) that in this case all $\alpha$-analytic solutions are of the form $v(t^{\alpha})$ where $v$ is an entire function of finite order of the growth. Finally, Theorem 4 gives sharp values for the order of the growth of $v$ under natural conditions on the coefficients. This improves the mentioned Kochubei’s result as a special case (see the corollary in Section 4). Auxiliary results are given in Section 2.
We use the notation $a\lesssim b$ if there exists a constant $C>0$ such that $a\leq Cb$. Similarly, $a\gtrsim b$ is understood in an analogous manner. If $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$ and say that $a$ and $b$ are comparable. Additionally, $a(t)\sim b(t)$ means that the quotient $a(t)/b(t)$ approaches one as $t$ tends its limit. By $[x]$ we denote the entire part of $x\in\mathbb{R}$.
@@ 1.2 Fractional integrals and derivatives
$I^{\alpha}x^{\beta-1}=\frac{\Gamma(\beta)}{\Gamma(\beta+\alpha)}x^{\beta+\alpha-1},\;D^{\alpha}x^{\beta-1}=\frac{\Gamma(\beta)}{\Gamma(\beta-\alpha)}x^{\beta-\alpha-1},\quad\alpha,\beta>0,\alpha\neq\beta,$ (1.2)
$D^{\alpha}1=\frac{1}{\Gamma(1-\alpha)}x^{-\alpha},\;D^{\alpha}x^{\alpha-j}=0,\quad\alpha>0,j\in\{1,2,\dots,[\alpha]+1\}.$ (1.3)
@@
$=D^{\alpha}\varphi(x)-\sum_{k=0}^{n-1}\frac{\varphi^{(k)}(0)}{\Gamma(k+1-\alpha)}x^{k-\alpha},\quad n-1<\alpha\leq n,$ (1.4)
@@
Let $H_{\alpha}(R)$, $0<R\leq\infty$, $\alpha>0$, denote the class of $\alpha$-analytic functions on $(0,R^{1/\alpha})$, that is the functions $u$ represented in the form $u(t)=\sum_{m=0}^{\infty}u_{m}t^{\alpha m}$, $\alpha>0$, $0\leq t^{\alpha}<R\leq\infty$. Direct computation yields
$(D^{\alpha}u)(t)=\sum_{m=0}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-1)\alpha+1)}t^{\alpha(m-1)},\quad 0<t^{\alpha}<R,$ (1.5)
@@
$=\sum_{m=1}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-1)\alpha+1)}t^{\alpha(m-1)}.$ (1.6)
@@
$(\mathbb{D}_{\alpha}^{j}u)(t)=\sum_{m=j}^{\infty}u_{m}\frac{\Gamma(m\alpha+1)}{\Gamma((m-j)\alpha+1)}t^{\alpha(m-j)}=\sum_{m=0}^{\infty}u_{m+j}\frac{\Gamma((m+j)\alpha+1)}{\Gamma(m\alpha+1)}t^{\alpha m}.$ (1.7)
@@
If $u\in H_{\alpha}(R)$, then $D^{\alpha}$ is associative and commutative providied the conditions $u_{m}=0$ for $m<\frac{\gamma+\beta}{2}$, $\gamma,\beta>0$, i.e.
@@
In particular, $\left(D^{\alpha}\right)^{j}=D^{j\alpha}$ if $u_{m}=0$ for $m<j$.
@@
We write $u(t)=\sum_{m=0}^{\infty}u_{m}t^{\alpha m}$, $\alpha>0$, $0<t^{\alpha}<R\leq\infty$. Since the power series is uniformly convergent on every segment $0,r^{\alpha}]\subset[0,R)$ (cf. [19, Theorem 3, Sec. IV.6]), we can integrate and differentiate it under the sum sign at every point of $0,R^{1/\alpha})$. Then using the fact that $u_{m}=0$ for $m<(\gamma+\beta)/\alpha$ we obtain
@@
$=\sum_{m=\left[\frac{\gamma+\beta}{\alpha}\right]}^{\infty}u_{m}\frac{\Gamma(\alpha m+1)}{\Gamma(\alpha m+1-\gamma)}\frac{\Gamma(\alpha m-\gamma+1)}{\Gamma(\alpha m+1-\gamma-\beta)}t^{\alpha m-\gamma-\beta}=D^{\gamma+\beta}u(t).$ (1.8)
@@
$f(z)=\sum\limits_{n=0}^{\infty}a_{n}z^{n},\quad z=re^{i\theta}$ (1.9)
@@
$(\mathbb{D}^{\alpha}u)(t)=\sum_{n=1}^{\infty}a_{n}\frac{\Gamma(n\alpha+1)}{\Gamma(n\alpha+1-\alpha)}t^{\alpha(n-1)}=(\mathcal{D}^{\alpha}f)(t^{\alpha}),$ (1.10)
@@
is the so-called Djrbashian-Gelfond-Leontiev operator ([15], [16], [20]), a special case of the Gelfond-Leontiev generalized differential operator corresponding to the Mittag-Leffler function $E_{\alpha}(z)=\sum_{k=0}^{\infty}\frac{z^{k}}{\Gamma(k\alpha+1)}$, $\alpha>0$.
The corresponding integral operator, right inverse to $\mathcal{D}_{\alpha}$, can be written as
$(\mathcal{I}_{\alpha}f)(z)=\sum_{n=0}^{\infty}a_{n}\frac{\Gamma(n\alpha+1)}{\Gamma(n\alpha+1-\alpha)}z^{n+1},$
@@
$\mathcal{I}_{\alpha}f(z)=\frac{z}{\Gamma(\alpha)}\int_{0}^{1}(1-t)^{\alpha-1}f(zt^{\alpha})\,dt.$
It follows from (1.10) that ([14, Section 2.5], [20, Sec. 18.2, 22.3])
$\mathcal{D}_{\alpha}=Q\circ\mathbb{D}_{\alpha}\circ Q^{-1},$
where $Q$ is the substitution operator $z\mapsto z^{\frac{1}{\alpha}}$, $Q^{-1}$ is its inverse, for a corresponding branch of a multivalued power function chosen on the segment $[0,z]$. This note allows us to apply the approach used in [7] to derive Wiman-Valiron type results for the operator $\mathcal{D}_{\alpha}$.
@@ 1.3 Wiman-Valiron theory
For $r\in[0,+\infty)$ and an entire function $f$ of the form (1.9) we denote $M(r,f)=\max\{|f(z)|:|z|=r\}$. We define the maximal term as $\mu(r,f)=\max\{|a_{n}|r^{n}:n\geq 0\}$ and the central index of the series as $\nu(r,f)=\max\{n\geq 0:|a_{n}|r^{n}=\mu(r,f)\}$. Note that $\nu(r,f)$ is non-decreasing, and an entire function $f$ is transcendental if and only if $\nu(r,f)\to+\infty$ as $r\to+\infty$.
@@
Let $V$ be the class of positive continuous nondecreasing functions $v$ on $0,+\infty)$ such that $\frac{x^{2}}{v(x)\ln v(x)}$ increases to $\infty$ on $x\in[x_{0};+\infty)$, $x_{0}>0$, and $\int\limits_{0}^{+\infty}\frac{dx}{v(x)}<+\infty$. For example, the functions $v_{0}(x)=x^{\beta}$ $(x\geq 0)$, $1<\beta<2$, and $v_{1}(x)=x\ln^{\alpha+1}x$, $(x\geq e)$, $\alpha\in(0,1)$, belong to $V$.
@@
$\mathcal{D}_{\alpha}^{j}f(z)=(\nu\alpha)^{j\alpha}\left(\frac{f(z)}{z^{j}}+O\left(\frac{\varkappa(\nu)}{\nu}\right)\frac{M(\rho,f)}{\rho^{j}}\right).$ (1.12)
@@
$M(\rho,\mathcal{D}_{\alpha}^{j}f(z))=\frac{(\nu\alpha)^{j\alpha}}{\rho^{j}}\left\{1+O\left(\frac{\varkappa(\nu)}{\nu}\right)\right\}M(\rho,f)=(1+o(1))\frac{(\nu\alpha)^{j\alpha}}{r^{j}}M(r,f)$ (1.13)